OpenStudy (anonymous):

Use number properties to simplify the following expression. -5 + (5 + 3)

OpenStudy (anonymous):

-5 +5 +3 = 3

OpenStudy (alesia/bailey):

number properties?

OpenStudy (anonymous):

a + (b +c) = (a + b) +c

OpenStudy (alesia/bailey):

so would that just be -5+(5+3)=(-5+5)+3

OpenStudy (anonymous):

Sorry for the delay. Yes that is correct.

OpenStudy (anonymous):

3 correct?

OpenStudy (anonymous):

3 is the answer.